# SyntheT2C: Generating Synthetic Data for Fine-Tuning Large Language Models on the Text2Cypher Task

### v1

@@ 4.2 Evaluation dataset and metrics
We utilize a dataset comprising 300 manually annotated and verified samples to evaluate our experiments. This dataset includes 150 questions annotated based on the Hetionet and LHY databases, respectively. Take Hetionet as an example, for every category among the 12 categories generated in Section 3.2.1, we employ GPT-3.5-Turbo to generate 10 new questions, forming 120 “in-domain” questions. Additionally, we introduce 3 unseen categories and generate 10 new questions for each new category, totaling 30 “out-of-domain” questions. For each of the 300 questions, the authors write a ground-truth Cypher query, which is then executed against the two databases to get the ground-truth execution results.
@@ 4.2.1 Evaluation of Cypher quality
The backbone LLMs, both pre-SFT and post-SFT, are tasked with generating Cyphers for the 300 questions in the evaluation dataset. Using GPT-4o OpenAI (2023), we determine the superior Cypher from the two provided versions. For each pair of Cyphers, we conduct two evaluations by varying the order of presenting the Cypher queries in the prompt to mitigate order-induced bias. If evaluations of both orders yield identical results, this judgment is accepted as the final outcome; otherwise, it is deemed a “Draw”.
@@ 4.3 Experiment setup
4.3.1 Cypher LLMs
Extensive experiments are conducted with four LLMs, including open/closed-source models. For open-source models, we evaluate Llama3, Qwen2 and InternLM2. For closed-sources model, we test GPT. The exact versions of the backbone LLMs are listed in Appendix E.
@@ 4.4 Supervised Fine-Tuning experiments
The backbone LLMs are fine-tuned with the MedT2C dataset, comprising 750 samples generated with the two pipelines and two Neo4j databases, totaling 3000 samples. The MedT2C dataset contains high-quality Question-Cypher pairs that passed all the automatic validations as well as the manual validation. In Appendix F we report the passing rate of each validator as a guide for further improvement of MedT2C’s data quality.
@@
In Figure 5, the win rates are calculated in comparison with the ground-truth Cyphers. We further conduct an experiment to compare directly the $c_{1}$ and $c_{2}$ generated with the same LLM with GPT-4o, using the prompt presented in Appendix G. The comparison results are shown in Figure 6. From these results, we can conclude that while the improvement may appear minor when comparing against the ground-truth Cyphers, a visible enhancement in Cypher quality is evident when comparing to the Cyphers generated by the pre-SFT model. We explain this difference as follows: the human annotations have a far higher quality than the Cyphers generated by vanilla LLMs. Therefore, even though the LLMs are enhanced after SFT, their output is still inferior to the human-annotated Cyphers, which is why the evaluation results in Figure 5 seem largely unchanged.
@@ 4.6 Ablation experiments
As presented in Table 1, when we use only the data generated by the template-filling pipeline to SFT the Llama3 model, the model’s performance actually declined after SFT. This can be attributed to the design of the template-filling pipeline, which emphasizes generating syntactically complex Cypher queries. When SFT is performed using only this data, the backbone LLM tends to produce unnecessarily complicated Cypher queries (e.g., breaking one query into two and then merging them). While this ability to write more complex Cypher queries is not directly reflected in the evaluation metrics, as the “hard questions” requiring advanced syntactic knowledge constitute only a small portion of the evaluation dataset, such data can enhance the LLM’s generalization capacity when combined with data from the LLM-based prompting pipeline.
@@
As shown in Table 2, each individual validator contributes some improvement, either in terms of Cypher quality or the accuracy of the execution results. Notably, the combination of all five validators yields the most significant increase in performance. This can be attributed to the validators’ collective ability to mitigate the majority of the bugs in the SFT dataset, thereby enhancing the overall quality of the generated Cypher queries.
@@ 5 Limitations
The primary limitation of our work is the challenge in writing the templates. While the templates are designed to be independent from the base Neo4j databases, some adaptation work is still necessary when applying them to new Neo4j databases. Additionally, writing new templates is time-consuming, making the expansion of the current template library difficult. Furthermore, a significant number of the generated Cypher queries are directly filtered out during the construction of MedT2C, resulting in a waste of resources. Developing methods to quickly fix Cyphers that do not satisfy all the validation criteria, instead of simply regenerating more Cyphers, could help reduce the carbon footprint of SyntheT2C.
@@ 6 Potential risks
Even though SyntheT2C is designed to automatically generate synthetic datasets, its usage requires close monitoring and manual validation to prevent the inadvertent inclusion of private or sensitive information. Additionally, the post-SFT LLM should be used with caution. Despite improvements in their Cypher generation performance, there remains a slight risk of producing embedded Cyphers that could lead to issues such as Out-of-Memory.
@@ B.3 Prompt to merge categories of questions
Appendix C Prompts used in automatic validators
C.1 Prompt of Semantic Validator
@@
C.2 Prompt of Coherence Validator
@@
Appendix D Important statistics of the LHY and the Hetionet databases
@@
Appendix E Exact versions of the backbone LLMs
@@
Appendix F Passing rate of MedT2C for each automatic validator
@@
Appendix G Prompts used for Cypher quality evaluation
We use GPT-4o to judge the quality of two versions of Cypher queries corresponding to the same set of questions written in natural language. The prompt used for this part is shown in Figure 15. We provide different aspects of evaluation and ask GPT-4o to give detailed reasons when evaluating, because these techniques bring more accurate evaluation results in practice.

### v2

@@ 4.2 Evaluation dataset and metrics
We utilize a dataset comprising 300 manually annotated and verified samples to evaluate our experiments. This dataset includes 150 questions annotated based on the Hetionet and LHY databases, respectively. Take Hetionet as an example, for every category among the 12 categories generated in Section 3.2.1, we employ GPT-3.5-Turbo to generate 10 new questions, forming 120 “in-domain” questions. Additionally, we introduce 3 unseen categories and generate 10 new questions for each new category, totaling 30 “out-of-domain” questions. For each of the 300 questions, the authors write a ground-truth Cypher query and test them manually to get the ground-truth execution results.
@@ 4.2.1 Evaluation of Cypher quality
The backbone LLMs, both pre-SFT and post-SFT, are tasked with generating Cyphers for the 300 questions in the evaluation dataset. Using GPT-4o (OpenAI, 2023), we determine the superior Cypher from the two provided versions. For each pair of Cyphers, we conduct two evaluations by varying the order of presenting the Cypher queries in the prompt to mitigate order-induced bias. If evaluations of both orders yield identical results, this judgment is accepted as the final outcome; otherwise, it is deemed a “Draw”.
@@ 4.3 Experiment setup
4.3.1 Cypher LLMs (baselines)
Extensive experiments are conducted with four LLMs, including open/closed-source models. For open-source models, we evaluate Llama3, Qwen2 and InternLM2. For closed-sources model, we test GPT. The exact versions of the backbone LLMs are listed in Appendix F. Our experiments focus on the performance boost brought by the SFT, therefore, the pre-SFT vanilla LLMs are used as the natural baselines. Empirically, the vanilla LLMs of the parameter size around 7B perform poorly on Cypher writing task. Only about 15%-20% of the few-shot generated Cyphers are executable, and the rate of semantically correct Cyphers is even lower. Thus, the few-shot prompting has become a default choice for 7B-level LLMs on Cypher writing task, and all the prompts we used in the experiments are 2-shot prompts.
We selected these 7B level models for our experiments to highlight the effectiveness of our method to enhance the smaller models with the synthetic data constructed with larger models.
On top of this, we also report the performance of GPT-4 as a reference: when tested on the Evaluation dataset, GPT-4 achieves the averaged Execution Result Accuracy of 49.07% (zero-shot) and 50.42% (2-shot).
@@ 4.4 Supervised Fine-Tuning experiments
The backbone LLMs are fine-tuned with the MedT2C dataset. MedT2C contains 3000 samples in total, with 750 samples generated from each combination of the two pipelines and the two Neo4j databases. The MedT2C dataset contains high-quality Question-Cypher pairs that passed all the automatic validations as well as the manual validation. In Appendix G we report the passing rate of each validator as a guide for further improvement of MedT2C’s data quality.
@@
In Figure 5, the win rates are calculated in comparison with the ground-truth Cyphers. We further conduct an experiment to compare directly the $c_{1}$ and $c_{2}$ generated with the same LLM with GPT-4o, using the prompt presented in Appendix H. The comparison results are shown in Figure 6. From these results, we can conclude that while the improvement may appear minor when comparing against the ground-truth Cyphers, a visible enhancement in Cypher quality is evident when comparing to the Cyphers generated by the pre-SFT model. We explain this difference as follows: the human annotations have a far higher quality than the Cyphers generated by vanilla LLMs. Therefore, even though the LLMs are enhanced after SFT, their output is still inferior to the human-annotated Cyphers, which is why the evaluation results in Figure 5 seem largely unchanged.
@@ 4.6 Ablation experiments
As presented in Table 1, when we use only the data generated by the template-filling pipeline to SFT the Llama3 model, the post-SFT performance actually declines. The reasons are: (1) template-filling pipeline generates rigid Cyphers which are easier to overfit; (2) when SFT with only template-filling data, LLMs tend to produce unnecessarily complicated Cypher queries (e.g., breaking one query into two and then merging them). From a global perspective, using data from both pipelines (marked as “All”) can enhance the LLM’s generalization capacity, as shown in comparison with the result obtained by fine-tuning only using the data generated with pipeline 1 (marked as "only LLM"). In other words, even though the template-filling data seems to have a negative impact, adding them to the fine-tuning dataset can further improve the overall Cypher writing performance.
@@
Table 2 shows that each individual validator contributes some improvement, either in terms of Cypher quality or the accuracy of the execution results. Notably, the combination of all five validators yields the most significant increase in performance. This can be attributed to the validators’ collective ability to mitigate the majority of the bugs in the SFT dataset, thereby enhancing the overall quality of the generated Cypher queries.
@@ 5 Limitations
The primary limitation of our work is the challenge in writing the templates. Besides, some adaptation work is necessary when applying them to new Neo4j databases (more details about adaptation to new database are discussed in Appendix I). Also, a portion of the generated Cypher queries are directly filtered out during the construction of MedT2C, which is not the most efficient solution. Finally, our current work focuses only on Cypher query language, similar method can be applied on other structured query languages in future research.
@@ 6 Potential risks
Though SyntheT2C is designed to automatically generate synthetic datasets, its usage requires close monitoring to prevent the inadvertent inclusion of private or sensitive information. Additionally, there is a slight residual risk for post-SFT LLMs of producing endless embedded Cyphers, which could potentially lead to issues such as Out-of-Memory.
@@ 7 Conclusion
Acknowledgements
We would like to express our gratitude to all the supervisors and colleagues at Shanghai Artificial Intelligence Laboratory for their invaluable insights, feedback, and support throughout the research process. We also thank the reviewers for their constructive comments, which greatly improved the quality of this paper.
@@ B.3 Prompt to merge categories of questions
Appendix C Details about constructing the templates
Our templates are based on the 60 templates introduced in Cypher Generator (Onofrei, 2024). These templates were originally written for an older version of Neo4j, so the authors first correct and update them to match the syntax of Neo4j 5.13. Additionally, the authors refer to the official Neo4j documentation to create 20 new templates that cover the new syntax and new features in Neo4j 5.13. All the 80 templates are tested by filling them with sampled values and executing them manually. After testing, all templates are confirmed to be able to generate executable Cypher queries.
Appendix D Prompts used in automatic validators
D.1 Prompt of Semantic Validator
@@
D.2 Prompt of Coherence Validator
@@
Appendix E Important statistics of the LHY and the Hetionet databases
@@
Appendix F Exact versions of the backbone LLMs
@@
Appendix G Passing rate of MedT2C for each automatic validator
@@
Appendix H Prompts used for Cypher quality evaluation
We use GPT-4o to judge the quality of two versions of Cypher queries corresponding to the same set of questions written in natural language. The prompt used for this part is shown in Figure 15. We provide different aspects of evaluation and ask GPT-4o to give detailed reasons when evaluating because these techniques bring more accurate evaluation results in practice.
Appendix I Adaptation of templates to new database
The challenge of adapting a template-based pipeline to unseen databases is a known issue as described in the section of Limitations. It is even a common drawback of all template-based methods. However, we have made efforts to mitigate this inconvenience. Our templates are written based on the work Cypher Generator (Onofrei, 2024) and enriched referring to the official documentation of Cypher Query Language, aiming to cover all syntax patterns of Cypher. When adapting to a new database, we comment out the templates involving non-existent data types (e.g., there is no DATE data in LHY database, so we comment out the DATE-related templates when applying on LHY database). Similarly, developers can comment or un-comment these templates when applying them to other databases, using these templates directly without rewriting from scratch. This method has proven to be effective when we adapted the pipeline to Hetionet and LHY databases during our experiments.